# Neural Dueling Bandits

### v1

@@ Abstract
Contextual dueling bandit is used to model the bandit problems, where a learner’s goal is to find the best arm for a given context using observed noisy preference feedback over the selected arms for the past contexts. However, existing algorithms assume the reward function is linear, which can be complex and non-linear in many real-life applications like online recommendations or ranking web search results. To overcome this challenge, we use a neural network to estimate the reward function using preference feedback for the previously selected arms. We propose upper confidence bound- and Thompson sampling-based algorithms with sub-linear regret guarantees that efficiently select arms in each round. We then extend our theoretical results to contextual bandit problems with binary feedback, which is in itself a non-trivial contribution. Experimental results on the problem instances derived from synthetic datasets corroborate our theoretical results.
@@ 1 Introduction
Contextual dueling bandits (or preference-based bandits) [1, 2, 3] is a sequential decision-making framework that is widely used to model the contextual bandit problems [4, 5, 6, 7, 8] in which a learner’s goal is to find an optimal arm by sequentially selecting a pair of arms (also refers as a duel) and then observing noisy preference feedback (i.e., one arm is preferred over another) for the selected arms. Contextual dueling bandits has many real-life applications, e.g., online recommendation, ranking web search, comparing two text responses generated from LLMs, and rating two restaurants/ movies, especially in the applications where it is easier to observe preference between two arms than knowing the absolute reward for the selected arm. The preference feedback between two arms [^1] For more than two arms, the preferences are assumed to follow the Plackett-Luce model [1, 9]. is often assumed to follow the Bradley-Terry-Luce (BTL) model [2, 10, 11] in which the probability of preferring an arm is proportional to the exponential of its reward.
@@
Since the number of contexts (e.g., users of online platforms) and arms (e.g., movies/search results to recommend) can be very large (or infinite), the reward of an arm is assumed to be parameterized by an unknown function, e.g., a linear function [1, 2, 3]. However, the reward function may not always be linear in practice. To overcome this challenge, this paper parameterizes the reward function via a non-linear function, which needs to be estimated using the available preference feedback for selected arms. To achieve this, we can estimate the non-linear function by using either a Gaussian processes [12, 13, 14] or a neural network [7, 8]. However, due to the limited expressive power of the Gaussian processes, it fails when optimizing highly complex functions. In contrast, neural networks (NNs) possess strong expressive power and can model highly complex functions [15, 16].
@@
In this paper, we first introduce the problem setting of neural dueling bandits, in which we use a neural network to model the unknown reward function in contextual dueling bandits. As compared to the existing work on neural contextual bandits [7, 8], we have to use cross-entropy loss due to binary preference feedback. We then propose the first two neural dueling bandit algorithms based on, respectively, upper confidence bound (UCB) [4, 5, 7, 17, 18, 19] and Thompson sampling (TS) [8, 14, 20, 21] (in Section 3.1). Under mild assumptions, we derive an upper bound on the estimation error of the difference between the reward values of any pair of arms (1), which is valid as long as the neural network is sufficiently wide. This result provides a theoretical assurance of the quality of our trained neural network using the preference feedback. Based on the theoretical guarantee on the estimation error, we derive upper bounds on the cumulative regret of both of our algorithms (2 and 3), which are sub-linear under some mild conditions. Our regret upper bounds lead to a number of interesting and novel insights (more details in Section 3.2).
@@
Interestingly, our theoretical results provide novel theoretical insights regarding the reinforcement learning with human feedback (RLHF) algorithm (Section 4). Specifically, our 1 naturally provides a theoretical guarantee on the quality of the learned reward model in terms of its accuracy in estimating the reward differences between pairs of responses. As a special case, we extend our results to neural contextual bandit problems with binary feedback in Section 5, which is itself of independent interest. Finally, we empirically validate the different performance aspects of our proposed algorithms in Section 6 using problem instances derived from synthetic datasets.
Related work.  Learning from pairwise or $K$-wise comparisons has been thoroughly explored in the literature. In the context of dueling bandits, the focus is on minimizing regret using preference feedback [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. We refer the readers to [35] for a detailed survey on dueling bandits. The closest work to ours is contextual dueling bandits [1, 2, 3, 36, 37], but they only consider the linear reward function, which we extend to the non-linear reward functions.
@@ 2 Problem Setting
Contextual dueling bandits.  We consider a contextual dueling bandit problem in which a learner selects two arms (also refers as a duel) for a given context and observes preference feedback over arms. The learner’s goal is to find the best arm for each context. Let $\mathcal{C}\subset\mathbb{R}^{d_{c}}$ be the context set and $\mathcal{A}\subset\mathbb{R}^{d_{a}}$ be finite arm set, where $d_{c}\geq 1$ and $d_{a}\geq 1$. At the beginning of round $t$, the environment generates a context $c_{t}\in\mathcal{C}$ and the learner selects two arms (i.e., $a_{t,1}$, and $a_{t,2}$) from the finite arm set $\mathcal{A}$. After selecting two arms, the learner observes stochastic preference feedback $y_{t}$ for the selected arms, where $y_{t}=1$ implies the arm $a_{t,1}$ is preferred over arm $a_{t,2}$ and $y_{t}=0$ otherwise. We assume that the preference feedback depends on an unknown non-linear reward function $f:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}$. For brevity, we denote the set of all context-arm feature vectors in the round $t$ by $\mathcal{X}_{t}\subset\mathbb{R}^{d}$ and use $x_{t}^{a}$ to represent the context-arm feature vector for context $c_{t}$ and an arm $a$.
Stochastic preference model.  We assume the preference has a Bernoulli distribution that follows the Bradley-Terry-Luce (BTL) model [10, 11], which is commonly used in the dueling bandits [1, 2, 3]. Under BTL preference model, the probability that the first selected arm $(x_{t,1})$ is preferred over the second selected arm $(x_{t,2})$ for the given reward function $f$ is given by
@@
where $x_{1}\succ x_{2}$ denotes that $x_{t,1}$ is preferred over $x_{t,2}$, $\mu(x)=1/(1+\mathrm{e}^{-x})$ [^2] Our results can be extended to other preference models like the Thurstone-Mosteller model and Exponential Noise as long as the stochastic transitivity holds [2]. is the logistic function, and $f(x_{t,i})$ is the latent reward of the $i$-th selected arm for the given context $c_{t}$. We need the following standard assumptions on function $\mu$ (also known as a link function in the literature [2, 19]):
@@
  • The link function $\mu:\mathbb{R}\rightarrow[0,1]$ is continuously differentiable and Lipschitz with constant $L_{\mu}$. For logistic function, $L_{\mu}\leq 1/4$.
Performance measure.  After selecting two arms, denoted by $x_{t,1}$ and $x_{t,2}$, in round $t$, the learner incurs an instantaneous regret. There are two common notions of instantaneous regret in the dueling bandits setting, i.e., average instantaneous regret: $r_{t}^{a}\doteq f(x_{t}^{\star})-\left({f(x_{t,1})+f(x_{t,2})}\right)/{2}$, and weak instantaneous regret: $r_{t}^{w}\doteq f(x_{t}^{\star})-\max\left\{f(x_{t,1}),f(x_{t,2})\right\}$, where $x_{t}^{\star}=\arg\!\max_{x\in\mathcal{X}_{t}}f(x)$ denotes the best arm for a given context that maximizes the value of the underlying reward function. After observing preference feedback for $T$ pairs of arms, the cumulative regret (or regret, in short) of a sequential policy is given by $\mathfrak{R}_{T}^{\tau}=\sum_{t=1}^{T}r_{t}^{\tau},$ where $\tau\doteq\{a,w\}$. Any good policy should have sub-linear regret, i.e., $\lim_{T\to\infty}{\mathfrak{R}_{T}^{\tau}}/T=0.$ A policy with a sub-linear regret implies that the policy will eventually find the best arm and recommend only the best arm in the duel for the given contexts.
@@ 3 Neural Dueling Bandits
Having a good reward function estimator is the key for any contextual bandit algorithm to achieve good performance, i.e., smaller regret. As the underlying reward function is complex and non-linear, we use fully connected neural networks [7, 8] to estimate the reward function only using the preference feedback. Using this estimated reward function, we propose two algorithms based on the upper confidence bound and Thomson sampling with sub-linear regret guarantees.
Reward function estimation using neural network.
To estimate the unknown reward function $f$, we use a fully connected neural network (NN) with depth $L\geq 2$, the width of hidden layer $m$, and ReLU activations as done in [7] and [8]. Let $h(x;\theta)$ represent the output of a full-connected neural network with parameters $\theta$ for context-arm feature vector $x$, which is defined as follows:
@@
Here $\theta_{0}$ represents the initial parameters of the NN, and we initialize $\theta_{0}$ following the standard practice of neural bandits [7, 8] (refer to Algorithm 1 in [8] for details). Here, minimizing the first term in the loss function (i.e., the term involving the summation from $t-1$ terms) corresponds to the maximum log likelihood estimate (MLE) of the parameters $\theta$. Next, we develop algorithms that use the trained NN with parameter $\theta_{t}$ to select the best arms (duel) for each context.
@@ 3.1 Neural dueling bandit algorithms
UCB-based algorithm.  Using upper confidence bound for dealing with the exploration-exploitation trade-off is common in many sequential decision-making problems [2, 7, 17]. We propose a UCB-based algorithm named NDB-UCB, which works as follows: At the beginning of the round $t$, the algorithm trains the NN using available observations. After receiving the context, it selects the first arm greedily (i.e., by maximizing the output of the trained NN with parameter $\theta_{t}$) as follows:
@@
Thompson sampling [3, 21] selects an arm according to its probability of being the best. Many works [3, 14, 21, 38] have shown that TS is empirically superior than to its counterpart UCB-based bandit algorithms. Therefore, in addition, we also propose another algorithm based on TS named NDB-TS, which works similarly to NDB-UCB except that the second arm $x_{t,2}$ is selected differently. To select the second arm $x_{t,2}$, for every arm $x\in\mathcal{X}_{t}$, it firstly samples a reward $r_{t}(x)\sim\mathcal{N}\left(h(x;\theta_{t})-h(x_{t,1};\theta_{t}),\nu_{T}^{2}\sigma_{t-1}^{2}(x,x_{t,1})\right)$ and then selects the second arm as $x_{t,2}={\arg\max}_{x\in\mathcal{X}_{t}}{r}_{t}(x)$.
@@ 3.2 Regret analysis
Let $K$ denote the finite number of available arms in each round, $\mathbf{H}$ denote the NTK matrix for all $T\times K$ context-arm feature vectors in the $T$ rounds, and $h=\left(f(x_{1}^{1}),\ldots,f(x_{T}^{K})\right)$. The NTK matrix $\mathbf{H}$ definition is adapted to our setting from Definition 4.1 of [7]. We now introduce the assumptions needed for our regret analysis, all of which are standard assumptions in neural bandits [7, 8].
@@
  • all context-arm feature vectors satisfy $\left\|x\right\|_{2}=1$ and $[x]_{j}=[x]_{j+d/2}$, $\forall x\in\mathcal{X}_{t},\forall t\in[T]$.
The last assumption in 2 above, together with the way we initialize $\theta_{0}$ (i.e., following standard practice in neural bandits [7, 8]), ensures that $h(x;\theta_{0})=0,\forall x\in\mathcal{X}_{t},\forall t\in[T]$.
@@
Compared to the previous works on neural bandits, our definition of $\widetilde{d}$ features extra dependencies on $\kappa_{\mu}$. Moreover, our $\mathbf{H}^{\prime}$ contains $T\times K\times(K-1)$ contexts, which is more than the $T\times K$ contexts of [7] and [8].333 The effective dimension in [7] and [8] is defined using $\mathbf{H}$: $\widetilde{d}^{\prime}=\log\det\left(\mathbf{H}/\lambda+\mathbf{I}\right)/\log(1+TK/\lambda)$. However, it is of the same order (up to log factors) as $\log\det\left(\widetilde{\mathbf{H}}/\lambda+\mathbf{I}\right)$, with $\widetilde{\mathbf{H}}\doteq\sum_{s=1}^{T}\sum^{K}_{i=1}g(x_{s,i};\theta_{0})g(x_{s,i};\theta_{0})^{\top}/m$ (see Lemma B.7 of [8]). Hence, our $\widetilde{d}$ is expected to be generally larger than their standard effective dimension.
Note that placing an assumption on $\widetilde{d}$ above is analogous to the assumption on the eigenvalue of the matrix $M_{t}$ in the work on linear dueling bandits [2]. For example, in order for our final regret bound to be sub-linear, we only need to assume that $\widetilde{d}=\widetilde{O}(\sqrt{T})$, which is analogous to the assumption from [2]: $\sum^{T}_{t=\tau+1}\lambda_{\min}^{-1/2}(M_{t})\leq c\sqrt{T}$.
A key step in our proof is that minimizing the loss function Eq. 1 allows us to achieve the following:
$\frac{1}{m}\sum_{s=1}^{t-1}\left(\mu\left(h(x_{s,1};\theta_{t})-h(x_{s,2};\theta_{t})\right)-y_{s}\right)(g(x_{s,1};\theta_{t})-g(x_{s,2};\theta_{t}))+\lambda(\theta_{t}-\theta_{0})=0.$ (5)
We use the above fact to prove our following result, which is equivalent to the confidence ellipsoid results used in the existing bandit algorithms [19].
@@
$\mathfrak{R}_{T}\leq\frac{3}{\sqrt{2}}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\sqrt{c_{0}T\widetilde{d}}+1=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
@@
Under the same conditions as those in 2, then with probability of at least $1-\delta$, we have
@@
Note that in terms of asymptotic dependencies (ignoring the log factors), our UCB- and TS- algorithms have the same growth rates.
@@
As we have discussed above, if we place an assumption on the effective dimension $\widetilde{d}=\widetilde{O}(\sqrt{T})$ (which is analogous to the assumption on the minimum eigenvalue from [2]), then the regret upper bounds for both NDB-UCB and NDB-TS are sub-linear. The dependence of our regret bounds on $\frac{1}{\kappa_{\mu}}$ and $L_{\mu}$ (i.e., the parameters of the link function defined in 1) are consistent with the previous works on generalized linear bandits [19] and linear dueling bandits [2].

### v2

@@ Abstract
Contextual dueling bandit is used to model the bandit problems, where a learner’s goal is to find the best arm for a given context using observed noisy human preference feedback over the selected arms for the past contexts. However, existing algorithms assume the reward function is linear, which can be complex and non-linear in many real-life applications like online recommendations or ranking web search results. To overcome this challenge, we use a neural network to estimate the reward function using preference feedback for the previously selected arms. We propose upper confidence bound- and Thompson sampling-based algorithms with sub-linear regret guarantees that efficiently select arms in each round. We also extend our theoretical results to contextual bandit problems with binary feedback, which is in itself a non-trivial contribution. Experimental results on the problem instances derived from synthetic datasets corroborate our theoretical results.
@@ 1 Introduction
Contextual dueling bandits (or preference-based bandits) (Saha, 2021; Bengs et al., 2022; Li et al., 2024) is a sequential decision-making framework that is widely used to model the contextual bandit problems (Li et al., 2010; Chu et al., 2011; Krause & Ong, 2011; Zhou et al., 2020; Zhang et al., 2021) in which a learner’s goal is to find an optimal arm by sequentially selecting a pair of arms (also refers as a duel) and then observing noisy human preference feedback (i.e., one arm is preferred over another) for the selected arms. Contextual dueling bandits has many real-life applications, such as online recommendation, ranking web search, fine-tuning large language models, and rating two restaurants or movies, especially in the applications where it is easier to observe human preference between two options (arms) than knowing the absolute reward for the selected option (arm). The preference feedback between two arms [^1] For more than two arms, the preferences are assumed to follow the Plackett-Luce model (Saha, 2021; Soufiani et al., 2014). is often assumed to follow the Bradley-Terry-Luce (BTL) model (Hunter, 2004; Luce, 2005; Saha, 2021; Bengs et al., 2022; Li et al., 2024) in which the probability of preferring an arm is proportional to the exponential of its reward.
@@
Since the number of contexts (e.g., users of online platforms) and arms (e.g., movies/search results to recommend) can be very large (or infinite), the reward of an arm is assumed to be parameterized by an unknown function, e.g., a linear function (Saha, 2021; Bengs et al., 2022; Li et al., 2024). However, the reward function may not always be linear in practice. To overcome this challenge, this paper parameterizes the reward function via a non-linear function, which needs to be estimated using the available preference feedback for selected arms. To achieve this, we can estimate the non-linear function by using either a Gaussian processes (Williams & Rasmussen, 2006; Srinivas et al., 2010; Chowdhury & Gopalan, 2017) or a neural network (Zhou et al., 2020; Zhang et al., 2021). However, due to the limited expressive power of the Gaussian processes, it fails when optimizing highly complex functions. In contrast, neural networks (NNs) possess strong expressive power and can model highly complex functions (Dai et al., 2023; Lin et al., 2023).
In this paper, we first introduce the problem setting of neural dueling bandits, in which we use a neural network to model the unknown reward function in contextual dueling bandits. As compared to the existing work on neural contextual bandits (Zhou et al., 2020; Zhang et al., 2021), we have to use cross-entropy loss as an objective function for training the neural network to estimate the unknown non-linear reward function due to the preference feedback (i.e., 00/$1$). We next propose two neural dueling bandit algorithms based on, respectively, upper confidence bound (UCB) (Auer et al., 2002; Li et al., 2010; Chu et al., 2011; Abbasi-Yadkori et al., 2011; Li et al., 2017; Zhou et al., 2020) and Thompson sampling (TS) (Krause & Ong, 2011; Agrawal & Goyal, 2013a; b; Chowdhury & Gopalan, 2017; Zhang et al., 2021) (more details are in Section 3.2). Note that the existing contextual dueling bandit works (Saha, 2021; Bengs et al., 2022; Li et al., 2024) use different ways to select the pair of arms (more details are given in Section C.2), hence leading to different arm-selection strategies as compared to our work. Due to the differences in arm-selection strategy and reward function (which is non-linear and estimated using NNs), our regret analysis is also completely different.
@@
Furthermore, existing neural contextual bandit works use root mean square error (RMSE) as an objective function for training the neural networks due to the assumption of real-valued reward or feedback. Therefore, their key results, especially the confidence ellipsoid, only hold for RMSE and can not be straightforwardly extended to our setting, which uses cross-entropy loss. Nevertheless, we derive an upper bound on the estimation error (i.e., represented as a confidence ellipsoid) of the difference between the reward values of any pair of arms (1) predicted by the trained neural network, which is valid as long as the neural network is sufficiently wide. This result provides a theoretical assurance of the quality of our trained neural network using the preference feedback to minimize the cross-entropy loss. Based on the theoretical guarantee on the estimation error, we derive upper bounds on the cumulative regret of both of our algorithms (2 and 3), which are sub-linear under some mild conditions. Our regret upper bounds lead to several interesting and novel insights (more details are given in Section 3.3).
@@
As a special case, we extend our results to neural contextual bandit problems with binary feedback in Appendix D, which is itself of independent interest. Interestingly, our theoretical results also provide novel theoretical insights regarding the reinforcement learning with human feedback (RLHF) algorithm (Appendix E). Specifically, our 1 naturally provides a theoretical guarantee on the quality of the learned reward model in terms of its accuracy in estimating the reward differences between pairs of responses. Finally, we empirically validate the different performance aspects of our proposed algorithms in Section 5 using problem instances derived from synthetic datasets.
@@ 2 Problem Setting
Contextual dueling bandits.  We consider a contextual dueling bandit problem in which a learner selects two arms (also refers as a duel) for a given context and observes preference feedback over selected arms. The learner’s goal is to find the best arm for each context. Our problem differs from standard contextual bandits in which a learner selects a single arm and observes an absolute numerical reward for that arm. Let $\mathcal{C}\subset\mathbb{R}^{d_{c}}$ be the context set and $\mathcal{A}\subset\mathbb{R}^{d_{a}}$ be finite arm set, where $d_{c}\geq 1$ and $d_{a}\geq 1$. At the beginning of round $t$, the environment generates a context $c_{t}\in\mathcal{C}$ and the learner selects two arms (i.e., $a_{t,1}$, and $a_{t,2}$) from the finite arm set $\mathcal{A}$. After selecting two arms, the learner observes stochastic preference feedback $y_{t}$ for the selected arms, where $y_{t}=1$ implies the arm $a_{t,1}$ is preferred over arm $a_{t,2}$ and $y_{t}=0$ otherwise. We assume that the preference feedback depends on an unknown non-linear reward function $f:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}$. For brevity, we denote the set of all context-arm feature vectors in the round $t$ by $\mathcal{X}_{t}$. We also use $\mathcal{X}$ to denote the set of all feature vectors: $\mathcal{X}_{t}\subset\mathcal{X},\forall t$ and $x_{t,a}$ to represent the context-arm feature vector for context $c_{t}$ and an arm $a$.
Stochastic preference model.  We assume the preference has a Bernoulli distribution that follows the Bradley-Terry-Luce (BTL) model (Hunter, 2004; Luce, 2005), which is commonly used in the dueling bandits (Saha, 2021; Bengs et al., 2022; Li et al., 2024). Under the BTL preference model, the probability that the first selected arm $(x_{t,1})$ is preferred over the second selected arm $(x_{t,2})$ for the the given context $c_{t}$ and latent reward function $f$ is given by
@@
where $x_{t,1}\succ x_{t,2}$ denotes that $x_{t,1}$ is preferred over $x_{t,2}$, $\mu(x)=1/(1+\mathrm{e}^{-x})$ is the sigmoid function and $f(x_{t,i})$ is the latent reward of the $i$-th selected arm. Our results hold for other preference models like the Thurstone-Mosteller model and Exponential Noise as long as the stochastic transitivity holds (Bengs et al., 2022). To generalize our results across preference models, we make the following assumptions on function $\mu$ (also known as a link function (Li et al., 2017; Bengs et al., 2022)):
@@
  • The link function $\mu:\mathbb{R}\rightarrow[0,1]$ is continuously differentiable and Lipschitz with constant $L_{\mu}$. For sigmoid function, $L_{\mu}\leq 1/4$.
Performance measure.  After selecting two arms, denoted by $x_{t,1}$ and $x_{t,2}$, in round $t$, the learner incurs an instantaneous regret. There are two common notions of instantaneous regret in the dueling bandits setting (Saha, 2021; Bengs et al., 2022; Li et al., 2024), i.e., average instantaneous regret: $r_{t}^{a}\doteq f(x_{t}^{\star})-\left({f(x_{t,1})+f(x_{t,2})}\right)/{2}$, and weak instantaneous regret: $r_{t}^{w}\doteq f(x_{t}^{\star})-\max\left\{f(x_{t,1}),f(x_{t,2})\right\}$, where $x_{t}^{\star}=\arg\!\max_{x\in\mathcal{X}_{t}}f(x)$ denotes the best arm for a given context that maximizes the value of the underlying reward function. After observing preference feedback for $T$ pairs of arms, the cumulative regret (or regret, in short) of a sequential policy is given by $\mathfrak{R}_{T}^{\tau}=\sum_{t=1}^{T}r_{t}^{\tau},$ where $\tau\doteq\{a,w\}$. Note that $\mathfrak{R}_{T}^{w}\leq\mathfrak{R}_{T}^{a}$. Any good policy should have sub-linear regret, i.e., $\lim_{T\to\infty}{\mathfrak{R}_{T}^{\tau}}/T=0.$ A policy with a sub-linear regret implies that the policy will eventually find the best arm and recommend only the best arm in the duel for the given contexts.
@@ 3 Neural Dueling Bandits
Having a good reward function estimator is the key for any contextual bandit algorithm to achieve good performance, i.e., smaller regret. As the underlying reward function is complex and non-linear, we use fully connected neural networks (Zhou et al., 2020; Zhang et al., 2021) to estimate the reward function only using the preference feedback. Using this estimated reward function, we propose two algorithms based on the UCB and TS with sub-linear regret guarantees.
3.1 Reward Function Estimation Using Neural Network
To estimate the unknown reward function $f$, we use a fully connected neural network (NN) with depth $L\geq 2$, the width of hidden layer $m$, and ReLU activations as done in Zhou et al. (2020) and Zhang et al. (2021). Let $h(x;\theta)$ represent the output of a full-connected neural network with parameters $\theta$ for context-arm feature vector $x$, which is defined as follows:
@@ 3.1 Reward Function Estimation Using Neural Network
Here $\theta_{0}$ represents the initial parameters of the NN, and we initialize $\theta_{0}$ following the standard practice of neural bandits (Zhou et al., 2020; Zhang et al., 2021) (refer to Algorithm 1 in Zhang et al. (2021) for details). Here, minimizing the first term in the loss function (i.e., the term involving the summation from $t-1$ terms) corresponds to the maximum log likelihood estimate (MLE) of the parameters $\theta$. Next, we develop algorithms that use the trained NN with parameter $\theta_{t}$ to select the best arms (duel) for each context.
@@ 3.2 Neural Dueling Bandit Algorithms
UCB-based algorithm.  Using upper confidence bound for dealing with the exploration-exploitation trade-off is common in many sequential decision-making problems (Bengs et al., 2022; Zhou et al., 2020; Auer et al., 2002). We propose a UCB-based algorithm named NDB-UCB, which works as follows: At the beginning of the round $t$, the algorithm trains the NN using available observations. After receiving the context, it selects the first arm greedily (i.e., by maximizing the output of the trained NN with parameter $\theta_{t}$) as follows:
@@
Thompson sampling (Li et al., 2024; Agrawal & Goyal, 2013b) selects an arm according to its probability of being the best. Many works (Li et al., 2024; Chowdhury & Gopalan, 2017; Agrawal & Goyal, 2013b; Chapelle & Li, 2011) have shown that TS is empirically superior than to its counterpart UCB-based bandit algorithms. Therefore, in addition, we also propose another algorithm based on TS named NDB-TS, which works similarly to NDB-UCB except that the second arm $x_{t,2}$ is selected differently. To select the second arm $x_{t,2}$, for every arm $x\in\mathcal{X}_{t}$, it firstly samples a reward $r_{t}(x)\sim\mathcal{N}\left(h(x;\theta_{t})-h(x_{t,1};\theta_{t}),\nu_{T}^{2}\sigma_{t-1}^{2}(x,x_{t,1})\right)$ and then selects the second arm as $x_{t,2}={\arg\max}_{x\in\mathcal{X}_{t}}{r}_{t}(x)$.
@@ 3.3 Regret analysis
Let $K$ denote the finite number of available arms in each round, $\mathbf{H}$ denote the NTK matrix for all $T\times K$ context-arm feature vectors in the $T$ rounds, and $h=\left(f(x_{1}^{1}),\ldots,f(x_{T}^{K})\right)$. The NTK matrix $\mathbf{H}$ definition is adapted to our setting from Definition 4.1 of Zhou et al. (2020). We denote the $j$-th element of the vector $x$ by $x_{j}$. We now introduce the assumptions needed for our regret analysis, all of which are standard assumptions in neural bandits Zhou et al. (2020); Zhang et al. (2021).
@@
  • all context-arm feature vectors satisfy $\left\|x\right\|_{2}=1$ and $x_{j}=x_{j+d/2}$, $\forall x\in\mathcal{X}_{t},\forall t\in[T]$.
The last assumption in 2 above, together with the way we initialize $\theta_{0}$ (i.e., following standard practice in neural bandits Zhou et al. (2020); Zhang et al. (2021)), ensures that $h(x;\theta_{0})=0,\forall x\in\mathcal{X}_{t},\forall t\in[T]$. The assumption of $x_{j}=x_{j+d/2}$ is a mild assumption and commonly used in the neural bandits literature Zhou et al. (2020); Zhang et al. (2021). This assumption is just for convenience in regret analysis: for any context $x$, $||x||=1$, we can always construct a new context $x^{\prime}=(x^{\top},x^{\top})^{\top}/\sqrt{2}$ that satisfies this assumption Zhou et al. (2020).
@@
Compared to the previous works on neural bandits, our definition of $\widetilde{d}$ features extra dependencies on $\kappa_{\mu}$. Moreover, our $\mathbf{H}^{\prime}$ contains $T\times K\times(K-1)$ contexts, which is more than the $T\times K$ contexts of Zhou et al. (2020) and Zhang et al. (2021).222 The effective dimension in Zhou et al. (2020) and Zhang et al. (2021) is defined using $\mathbf{H}$: $\widetilde{d}^{\prime}=\log\det\left(\mathbf{H}/\lambda+\mathbf{I}\right)/\log(1+TK/\lambda)$. However, it is of the same order (up to log factors) as $\log\det\left(\widetilde{\mathbf{H}}/\lambda+\mathbf{I}\right)$, with $\widetilde{\mathbf{H}}\doteq\sum_{s=1}^{T}\sum^{K}_{i=1}g(x_{s,i};\theta_{0})g(x_{s,i};\theta_{0})^{\top}/m$ (see Lemma B.7 of Zhang et al. (2021)). Hence, our $\widetilde{d}$ is expected to be larger than their standard effective dimension. It follows from the Determinant-Trace Inequality (Lemma 10 of Abbasi-Yadkori et al. (2011)), the determinant of a covariance matrix in directly proportional to the number of feature vectors. In our setting, the effective dimension $\widetilde{d}$ is $\log\det\left(\frac{\kappa_{\mu}}{\lambda}\mathbf{H}^{\prime}+\mathbf{I}\right)$ and $\mathbf{H}^{\prime}$ contains $T\times K\times(K-1)$ contexts (i.e., feature vectors), which is more than the $T\times K$ contexts in the standard neural bandits. Therefore, the effective dimension $\widetilde{d}$ in the neural dueling bandits is larger than that of the standard neural bandits.
Note that placing an assumption on $\widetilde{d}$ above is analogous to the assumption on the eigenvalue of the matrix $M_{t}$ in the work on linear dueling bandits Bengs et al. (2022). For example, in order for our final regret bound to be sub-linear, we only need to assume that $\widetilde{d}=\widetilde{o}(\sqrt{T})$, which is analogous to the assumption from Bengs et al. (2022): $\sum^{T}_{t=\tau+1}\lambda_{\min}^{-1/2}(M_{t})\leq c\sqrt{T}$.
We use the above fact to prove our following result, which is equivalent to the confidence ellipsoid results used in the existing bandit algorithms (Li et al., 2017).
@@
$\mathfrak{R}_{T}\leq\frac{3}{2}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\sqrt{T2c_{0}\widetilde{d}}+1=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
@@
Let $c_{1}$ and $c_{2}$ be two constants. Under the same conditions as those in 2, then with probability of at least $1-\delta$, we have
@@
Note that in terms of asymptotic dependencies (ignoring the log factors), our UCB- and TS- algorithms have the same growth rates. The upper regret bounds also hold for weak cumulative regret as it is upper-bounded by average cumulative regret.
@@
When we assume that the effective dimension $\widetilde{d}=\widetilde{o}(\sqrt{T})$ (which is analogous to the assumption on the minimum eigenvalue from Bengs et al. (2022)), then the regret upper bounds for both NDB-UCB and NDB-TS are sub-linear. The dependence of our regret bounds on $\frac{1}{\kappa_{\mu}}$ and $L_{\mu}$ (i.e., the parameters of the link function defined in 1) are consistent with the previous works on generalized linear bandits (Li et al., 2017) and linear dueling bandits Bengs et al. (2022).